OpenStudy (anonymous):

Find x3 − 2y2 − 3x3 + z4 if x = 3, y = 5, and z = −3.

OpenStudy (anonymous):

please give me medal

Mehek (mehek14):

do you have answer choices?

OpenStudy (anonymous):

I got -23

Mehek (mehek14):

\(3^3=27\\-2*5^2 = -50\\-3*3^3=-81\\-3^4=81\)

Mehek (mehek14):

\(27-(-50)=77\\77-(-81)=158\\158+81=239\)

OpenStudy (anonymous):

so is that the answer because I did it on a calculator and got -23

Mehek (mehek14):

yes that is what I got as the answer if you want I can double check

OpenStudy (anonymous):

that's good I was just wondering if you got what I got so I was double checking, but you can if you want.

Mehek (mehek14):

ok so \(3^3=27\\-2*5^2 = -50\\27-(-50)=77\)

Mehek (mehek14):

\(-3*3^3=-81\\77-(-81)=158\)

Mehek (mehek14):

\(-3^4=-81\\158+(-81)=77\)

Mehek (mehek14):

Well I go a different answer...

Mehek (mehek14):

yes it is 77

Mehek (mehek14):

I added in the last step instead of subtracting the first time I solved it

OpenStudy (anonymous):

here are the answer choices: a −185 b −77 c −23 d 85

Mehek (mehek14):

gotta check again

Mehek (mehek14):

I used my calculator and got -185

OpenStudy (anonymous):

okay thanks man hey can you fan me I am going to give you a medal can you give me one too

Mehek (mehek14):

sure

OpenStudy (anonymous):

oh and by the way I fanned you